Calculate the Gravimetric factor of the following. Show your Solution.


A. Al in Al2O3

B. Na in Na2C2O4

C. Cr2O3 in PbCrO4

D. K in K2PtCl6

E. Cl in AgCl


1
Expert's answer
2021-04-14T07:05:51-0400

Solution:

The gravimetric factor (GF), represents the weight of analyte per unit weight of precipitate. It is obtained from the ratio of the formula weight of the analyte to that of the precipitate, multiplied by the moles of analyte per mole of precipitate obtained from each mole of analyte, that is:



Thus:

A) Al in Al2O3

Analyte - Al; Precipitate - Al2O3

2Al → Al2O3

The molar mass of Al is 26.98154 g/mol.

The molar mass of Al2O3 is 101.96 g/mol.

Hence,

GF = (26.98154 g Al mol-1) / 101.96 g Al2O3 mol-1) × (2/1) = 0.52926 g Al / g Al2O3

GF = 0.52926 (Al in Al2O3)


B) Na in Na2C2O4

Analyte - Na; Precipitate - Na2C2O4

2Na → Na2C2O4

The molar mass of Na is 22.9897 g/mol.

The molar mass of Na2C2O4 is 133.9985 g/mol.

Hence,

GF = (22.9897 g Na mol-1) / 133.9985 g Na2C2O4 mol-1) × (2/1) = 0.34313 g Na / g Na2C2O4

GF = 0.34313 (Na in Na2C2O4)


C) Cr2O3 in PbCrO4

Analyte - Cr2O3; Precipitate - PbCrO4

Cr2O3 → 2PbCrO4

The molar mass of Cr2O3 is 151.99 g/mol.

The molar mass of PbCrO4 is 323.1937 g/mol.

Hence,

GF = (151.99 g Cr2O3 mol-1) / 323.1937 g PbCrO4 mol-1) × (1/2) = 0.23514 g Cr2O3 / g PbCrO4

GF = 0.23514 (Cr2O3 in PbCrO4)


D) K in K2PtCl6

Analyte - K; Precipitate - K2PtCl6

2K → K2PtCl6

The molar mass of K is 39.0983 g/mol.

The molar mass of K2PtCl6 is 485.99 g/mol.

Hence,

GF = (39.0983 g K mol-1) / 485.99 g K2PtCl6 mol-1) × (2/1) = 0.1609 g K / g K2PtCl6

GF = 0.16090 (K in K2PtCl6)


E) Cl in AgCl

Analyte - Cl; Precipitate - AgCl

Cl → AgCl

The molar mass of Cl is 35.453 g/mol.

The molar mass of AgCl is 143.32 g/mol.

Hence,

GF = (35.453 g Cl mol-1) / 143.32 g AgCl mol-1) × (1/1) = 0.24737 g Cl / g AgCl

GF = 0.24737 (Cl in AgCl)


Answers:

A) GF = GF = 0.52926 (Al in Al2O3);

B) GF = 0.34313 (Na in Na2C2O4);

C) GF = 0.23514 (Cr2O3 in PbCrO4);

D) GF = 0.16090 (K in K2PtCl6);

E) GF = 0.24737 (Cl in AgCl).